Find the slope of the line containing the given pair of points. If the line is undefined, state so. Please be sure to show all of your work. HELP!!! PLZ!!
(1, 2) and (9, 12)

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Slope, m, = diffy/diffx = (y2-y1)/(x2-x1)
(1,2) and (9,12)
x1 = 1, y1 = 2
x1 = 9, y2 = 12

Answer by stanbon(75887)

Find the slope of the line containing the given pair of points. If the line is undefined, state so. Please be sure to show all of your work. HELP!!! PLZ!!
(1, 2) and (9, 12)
----------------
slope = m = (12-2)/(9-1) = 10/8 = 5/4
=================
Cheers,
Stan H.